# Sequential Kernelized Stein Discrepancy

### v1

@@ 4.2 Composite null hypothesis
$g_{t}^{\theta}(x)=\frac{1}{\frac{1}{t-1}\sum_{i=1}^{t-1}M_{\theta}(X_{i})}\left(\frac{1}{t-1}\sum_{i=1}^{t-1}h_{\theta}(X_{i},x)\right)\quad\forall\theta\in\Theta.$
We note that, under the null hypothesis, there exists $\theta_{0}\in\Theta$ such that $Q=P_{\theta_{0}}$. Inspired by universal inference, we propose to consider the wealth process
@@ 5.1 A general approach
$\nabla_{x}\cdot\nabla_{\tilde{x}}k(x,\tilde{x})$ $=-3(1+\|x-\tilde{x}\|^{2})^{-\frac{5}{2}}\|x-\tilde{x}\|^{2}+d(1+\|x-\tilde{x}\|^{2})^{-\frac{3}{2}}.$
@@
$\langle s_{p}(\tilde{x}),\nabla_{x}k(x,\tilde{x})\rangle_{\mathbb{R}^{d}}+\langle s_{p}(x),\nabla_{\tilde{x}}k(x,\tilde{x})\rangle_{\mathbb{R}^{d}}$ $=\langle s_{p}(\tilde{x})-s_{p}(x),-\nabla_{x}k(x,\tilde{x})\rangle_{\mathbb{R}^{d}}$
@@
so it suffices to upper bound work with $s_{p}(\tilde{x})-s_{p}(x)$ and upper bound $\|s_{p}(\tilde{x})-s_{p}(x)\|$ by $\Gamma(\|x-\tilde{x}\|)+\Gamma^{\prime}(\tilde{x})$, where $\Gamma$ and $\Gamma^{\prime}$ are again appropriate functions. Note that (iii) has already been lower bounded. The choices of $\gamma,\gamma^{\prime},\Gamma,\Gamma^{\prime}$ will become clear in the next subsection, where we provide specific examples.
@@ 5.2 Specific examples of bound derivations
$\leq\left(|x-\tilde{x}|+|\tilde{x}-\theta|\right)|\tilde{x}-\theta|k(x,\tilde{x})$
@@
$(1+|x-\tilde{x}|^{2})^{-\frac{3}{2}}\langle s_{p}(\tilde{x})-s_{p}(x),x-\tilde{x}\rangle=-(1+|x-\tilde{x}|^{2})^{-\frac{3}{2}}|x-\tilde{x}|\in[-1,0].$
For a fixed $\tilde{x}$, it thus follows that
  • $|\langle s_{\theta}(x),s_{\theta}(\tilde{x})\rangle k(x,\tilde{x})|\leq|\tilde{x}-\theta|\{1+|\tilde{x}-\theta|\}$,
  • $\langle s_{\theta}(\tilde{x}),\nabla_{x}k(x,\tilde{x})\rangle+\langle s_{\theta}(x),\nabla_{\tilde{x}}k(x,\tilde{x})\rangle\in[-1,0]$,
  • $\nabla_{x}\cdot\nabla_{\tilde{x}}k(x,\tilde{x})\geq-2$.
@@
$\|\left(\theta_{1}(1-\tanh^{2}x_{1}),\theta_{2}(1-\tanh^{2}x_{2}),0\right)^{T}-x\|$ $\leq\|\left(\theta_{1}(\tanh^{2}\tilde{x}_{1}-\tanh^{2}x_{1}),\theta_{2}(\tanh^{2}\tilde{x}_{2}-\tanh^{2}x_{2}),0\right)^{T}\|$
$\quad+\|\tilde{x}-\left(\theta_{1}(1-\tanh^{2}\tilde{x}_{1}),\theta_{2}(1-\tanh^{2}\tilde{x}_{2}),0\right)^{T}\|+\|\tilde{x}-x\|$
@@
and so
$|\langle s_{\theta}(x),s_{\theta}(\tilde{x})\rangle k(x,\tilde{x})|$ $\leq\|s_{\theta}(x)\|\|s_{\theta}(\tilde{x})\|k(x,\tilde{x})$
@@
Secondly,
$\|s_{p}(\tilde{x})-s_{p}(x)\|\leq\|\theta\|+\|x-\tilde{x}\|,$
and so
$\left|(1+\|x-\tilde{x}\|^{2})^{-\frac{3}{2}}\langle s_{p}(\tilde{x})-s_{p}(x),x-\tilde{x}\rangle\right|$ $\leq(1+\|x-\tilde{x}\|^{2})^{-\frac{3}{2}}\|x-\tilde{x}\|\left(\|\theta\|+\|x-\tilde{x}\|\right)$
@@
For a fixed $\tilde{x}$, it thus follows that
  • $|\langle s_{\theta}(x),s_{\theta}(\tilde{x})\rangle k(x,\tilde{x})|\leq\left(\|\theta\|+\|s_{\theta}(\tilde{x})\|+1\right)\|s_{\theta}(\tilde{x})\|$,
  • $|\langle s_{\theta}(\tilde{x}),\nabla_{x}k(x,\tilde{x})\rangle+\langle s_{\theta}(x),\nabla_{\tilde{x}}k(x,\tilde{x})\rangle|\leq\|\theta\|+1$,
  • $\nabla_{x}\cdot\nabla_{\tilde{x}}k(x,\tilde{x})\geq 0$,
Hence it suffices to define $M_{\theta}(\tilde{x})=\left(\|\theta\|+\|s_{\theta}(\tilde{x})\|+1\right)\|s_{\theta}(\tilde{x})\|+\|\theta\|+1$. Note that $\mathbb{E}[M_{\theta}(X)]<\infty$ given that the distribution has Gaussian type tails, and so its first two moments are finite.
@@
where (i) is obtained given that $\phi(y)\in[-1,1]^{d_{h}}$ for all $y$. Thus
$|\langle s_{p}(x),s_{p}(\tilde{x})\rangle k(x,\tilde{x})|$ $\leq\|s_{p}(x)\|\|s_{p}(\tilde{x})\|k(x,\tilde{x})$
$\leq\left(\|s_{p}(\tilde{x})\|+\|\tilde{x}-x\|+\left\|\frac{B}{2}\phi(\frac{B^{T}x}{2}+c)-\frac{B}{2}\phi(\frac{B^{T}\tilde{x}}{2}+c)\right\|\right)\|s_{p}(\tilde{x})\|k(x,\tilde{x})$
$\leq\left(\|s_{p}(\tilde{x})\|+1+\|B\|_{op}\sqrt{d_{h}}\right)\|s_{p}(\tilde{x})\|.$
Secondly,
$\|s_{p}(\tilde{x})-s_{p}(x)\|\leq\|B\|_{op}\sqrt{d_{h}}+\|x-\tilde{x}\|,$
and so
$\left|(1+\|x-\tilde{x}\|^{2})^{-\frac{3}{2}}\langle s_{p}(\tilde{x})-s_{p}(x),x-\tilde{x}\rangle\right|$ $\leq(1+\|x-\tilde{x}\|^{2})^{-\frac{3}{2}}\|x-\tilde{x}\|\left(\|B\|_{op}\sqrt{d_{h}}+\|x-\tilde{x}\|\right)$
@@ 6 Experiments
We present here the empirical performance of the test. We consider the distributions introduced in the previous section, using the upper bounds derived therein. Gaussian distribution: We take $\theta=0$ under the null, and $\theta=1$ under the alternate. Intractable model: We take $\theta=(0,0)$ under the null, and $\theta=(1,1)$ under the alternate. Gaussian-Bernoulli Restricted Boltzmann Machine: We take $d_{h}=10$, and $d=50$. We sample from it using Gibbs sampling with a burn-in of 1000 iterations. Under the null, we take $b=0$, $c=0$, and matrix $B$ such that $B_{ij}$ is one if visible node $i$ is connected to hidden node $j$, and zero otherwise, with each hidden node connected to five visible nodes. We consider two different alternatives. For one of them, each entry of $B$ is shifted by $0.5$. For the other one, $b=1$.
Figure 1 exhibits the performance of the proposed test under the four alternatives considered, with $\alpha=0.05$. We emphasize the exponential growth of the wealth process. This behaviour is expected, in view of Theorem 2 and the fact that all these examples fulfil regularity conditions so that $\mathbb{E}_{H_{1}}[g^{*}(X)]>0$. Furthermore, we emphasize that aGRAPA empirically outperforms LBOW. This is due to the fact that $\lambda_{t+1}^{\text{aGRAPA}}>\lambda_{t+1}^{\text{LBOW}}$, so aGRAPA bets more aggressively. While not shown, we highlight that all the tests control the type-1 error at the desired level $0.05$. We defer to Appendix A.1 a discussion concerning the empirical performance of the proposed test and the fixed sample size kernel Stein discrepancy test, and to Appendix A.2 an empirical verification of the lower bound derived in Theorem 2.
Figure 1: Average logarithmic wealth alongside $95\%$ empirical confidence intervals under the alternative for (I) the Gaussian distribution, (II) the intractable model, (III) the restricted Boltzmann machine with shifted $B$, (IV) the restricted Boltzmann machine with bias $b=1$. We highlight the exponential growth of all wealth processes.
@@ 7 Conclusion
We have developed a sequential version of the kernel Stein discrepancy, which gives way to goodness-of-fit tests that can handle distributions with unknown normalizing constants and can be continuously monitored and adaptively stopped. We have done so by combining tools from testing by betting with RKHS theory, while avoiding assuming uniform boundedness of the Stein reproducing kernel. We have proved the validity of the test, as well as exponential growth of the wealth process under the alternative and mild regularity conditions. Our experiments have exhibited the empirical performance of the test in a variety of scenarios.
@@ Acknowledgements
DMT gratefully acknowledges that the project that gave rise to these results received the support of a fellowship from ‘la Caixa’ Foundation (ID 100010434). The fellowship code is LCF/BQ/EU22/11930075. AR was funded by NSF grant DMS-2310718.
@@ Appendix A Additional experiments
A.1 Comparison to batch setting
In Section 6, we exhibited the performance of the proposed, sequential test. We emphasize that the proposed test is anytime valid, counting with the subsequent substantial advantages that have been discussed in the main body of the work. Thus, it is expected to be outperformed in the batch setting by algorithms that are tailored to fixed sample sizes. Nonetheless, it is of interest to explore how the proposed test compares with the more classical kernel Stein discrepancy test. We present in this appendix the empirical performance of both the sequential and fixed sample size algorithms.

### v2

@@ 4.2 Composite Null Hypothesis
$g_{t}^{\theta}(x)=\frac{1}{\frac{1}{t-1}\sum_{i=1}^{t-1}M_{\theta}(X_{i})}\left(\frac{1}{t-1}\sum_{i=1}^{t-1}h_{\theta}(X_{i},x)\right)$
for $\theta\in\Theta$. We note that, under the null hypothesis, there exists $\theta_{0}\in\Theta$ such that $Q=P_{\theta_{0}}$. Inspired by universal inference, we propose to consider the wealth process
@@ 5.1 A General Approach
$\nabla_{x}\cdot\nabla_{\tilde{x}}k(x,\tilde{x})$ $=-3(1+\|x-\tilde{x}\|^{2})^{-\frac{5}{2}}\|x-\tilde{x}\|^{2}$
$\quad+d(1+\|x-\tilde{x}\|^{2})^{-\frac{3}{2}}.$
@@
$\langle s_{p}(\tilde{x}),\nabla_{x}k(x,\tilde{x})\rangle_{\mathbb{R}^{d}}+\langle s_{p}(x),\nabla_{\tilde{x}}k(x,\tilde{x})\rangle_{\mathbb{R}^{d}}$
$=$ $\langle s_{p}(\tilde{x})-s_{p}(x),-\nabla_{x}k(x,\tilde{x})\rangle_{\mathbb{R}^{d}}$
@@
so it suffices to upper bound work with $s_{p}(\tilde{x})-s_{p}(x)$ and upper bound $\|s_{p}(\tilde{x})-s_{p}(x)\|$ by $\Gamma(\|x-\tilde{x}\|)+\Gamma^{\prime}(\tilde{x})$, where $\Gamma$ and $\Gamma^{\prime}$ are again appropriate functions. Note that (iii) has already been lower bounded.
The choices of $\gamma,\gamma^{\prime},\Gamma,\Gamma^{\prime}$ will become clear in the next subsection, where we provide specific examples. We highlight that there should exist sensible choices as soon as $\|s_{p}(x)\|k(x,\tilde{x})$ and $\|s_{p}(x)-s_{p}(\tilde{x})\|\|\nabla k(x,\tilde{x})\|$ are both $O(1)$ for fixed $\tilde{x}$, i.e. $\|s_{p}(x)\|=O(\|x\|)$ in the case of the IMQ kernel. In addition to the cases considered in the next subsection, which include Gaussian distributions and restricted Boltzmann machines, a variety of models may fulfill this condition. In particular, exponential families with canonical parameters $\eta_{i}$ and canonical observations $T_{i}$ of the form $p_{\theta}(x)\propto\exp(\sum_{i\leq m}\eta_{i}(\theta)T_{i}(x))$ fall under this umbrella if $\|\nabla T_{i}(x)\|=O(\|x\|)$ for all $i\leq m$. For instance, exponential graphical models with linear interactions between nodes, which have found applications in protein signaling networks (see, e.g., Yang et al., (2015, Section 4.2.2.)) and spatial models (see, e.g., Besag, (1986, Section 4.2.2.) and Besag, (1974, Section 4.1.)) among others, attain such an assumption. Kernel exponential family models (Canu and Smola,, 2006) are another prominent example, which fulfill the condition as long as the gradients of the basis functions of a finite-rank approximation (as considered in Matsubara et al., (2022)) are $O(\|x\|)$ (which is precisely the case in Matsubara et al., (2022)).
@@ 5.2 Specific Examples of Bound Derivations
$\leq\left(|x-\tilde{x}|+|\tilde{x}-\theta|\right)|\tilde{x}-\theta|$
×$\quad\times k(x,\tilde{x})$
@@
$(1+|x-\tilde{x}|^{2})^{-\frac{3}{2}}\langle s_{p}(\tilde{x})-s_{p}(x),x-\tilde{x}\rangle$
$=$ $-(1+|x-\tilde{x}|^{2})^{-\frac{3}{2}}|x-\tilde{x}|\in[-1,0].$
For a fixed $\tilde{x}$, it thus follows that i) $|\langle s_{\theta}(x),s_{\theta}(\tilde{x})\rangle k(x,\tilde{x})|\leq|\tilde{x}-\theta|\{1+|\tilde{x}-\theta|\}$, ii) $\langle s_{\theta}(\tilde{x}),\nabla_{x}k(x,\tilde{x})\rangle+\langle s_{\theta}(x),\nabla_{\tilde{x}}k(x,\tilde{x})\rangle\in[-1,0]$, iii) $\nabla_{x}\cdot\nabla_{\tilde{x}}k(x,\tilde{x})\geq-2$.
@@
$\|\left(\theta_{1}(1-\tanh^{2}x_{1}),\theta_{2}(1-\tanh^{2}x_{2}),0\right)^{T}-x\|$
$\leq$ $\|\theta\|+\|\tilde{x}-x\|$
$+\|\tilde{x}-\left(\theta_{1}(1-\tanh^{2}\tilde{x}_{1}),\theta_{2}(1-\tanh^{2}\tilde{x}_{2}),0\right)^{T}\|$
@@
and so $|\langle s_{\theta}(x),s_{\theta}(\tilde{x})\rangle k(x,\tilde{x})|$ is upper bounded by
$\|s_{\theta}(x)\|\|s_{\theta}(\tilde{x})\|k(x,\tilde{x})$
@@
Secondly, $\|s_{p}(\tilde{x})-s_{p}(x)\|\leq\|\theta\|+\|x-\tilde{x}\|$, and so
$\left|(1+\|x-\tilde{x}\|^{2})^{-\frac{3}{2}}\langle s_{p}(\tilde{x})-s_{p}(x),x-\tilde{x}\rangle\right|$
$\leq$ $(1+\|x-\tilde{x}\|^{2})^{-\frac{3}{2}}\|x-\tilde{x}\|\left(\|\theta\|+\|x-\tilde{x}\|\right)$
@@
For a fixed $\tilde{x}$, it thus follows that i) $|\langle s_{\theta}(x),s_{\theta}(\tilde{x})\rangle k(x,\tilde{x})|\leq\left(\|\theta\|+\|s_{\theta}(\tilde{x})\|+1\right)\|s_{\theta}(\tilde{x})\|$, ii) $|\langle s_{\theta}(\tilde{x}),\nabla_{x}k(x,\tilde{x})\rangle+\langle s_{\theta}(x),\nabla_{\tilde{x}}k(x,\tilde{x})\rangle|\leq\|\theta\|+1$, iii) $\nabla_{x}\cdot\nabla_{\tilde{x}}k(x,\tilde{x})\geq 0$.
We thus define $M_{\theta}(\tilde{x})=\left(\|\theta\|+\|s_{\theta}(\tilde{x})\|+1\right)\|s_{\theta}(\tilde{x})\|+\|\theta\|+1$. Note that $\mathbb{E}[M_{\theta}(X)]<\infty$ given that the distribution has Gaussian type tails, and so its first two moments are finite.
@@
where (i) is obtained given that $\phi(y)\in[-1,1]^{d_{h}}$ for all $y$. Thus $|\langle s_{p}(x),s_{p}(\tilde{x})\rangle k(x,\tilde{x})|\leq\|s_{p}(x)\|\|s_{p}(\tilde{x})\|k(x,\tilde{x})$ is upper bounded by $\left(\|s_{p}(\tilde{x})\|+\|\tilde{x}-x\|+\left\|\frac{B}{2}\phi(\frac{B^{T}x}{2}+c)-\frac{B}{2}\phi(\frac{B^{T}\tilde{x}}{2}+c)\right\|\right)$$\times\|s_{p}(\tilde{x})\|k(x,\tilde{x})$, which is again upper bounded by
$\left(\|s_{p}(\tilde{x})\|+1+\|B\|_{op}\sqrt{d_{h}}\right)\|s_{p}(\tilde{x})\|.$
Secondly, $\|s_{p}(\tilde{x})-s_{p}(x)\|\leq\|B\|_{op}\sqrt{d_{h}}+\|x-\tilde{x}\|$, and so $\left|(1+\|x-\tilde{x}\|^{2})^{-\frac{3}{2}}\langle s_{p}(\tilde{x})-s_{p}(x),x-\tilde{x}\rangle\right|$ is upper bounded by
$(1+\|x-\tilde{x}\|^{2})^{-\frac{3}{2}}\|x-\tilde{x}\|\left(\|B\|_{op}\sqrt{d_{h}}+\|x-\tilde{x}\|\right)$
@@ 6 EXPERIMENTS
Figure 1: Average logarithmic wealth alongside $95\%$ empirical confidence intervals for $1000$ simulations under the alternative for (I) the Gaussian distribution, (II) the intractable model. We highlight the exponential growth of the wealth.
We consider the distributions introduced in the previous section, using the upper bounds derived therein. Gaussian distribution: We take $\theta_{0}=0$ under the null. Intractable model: We take $\theta_{0}=(0,0)$ under the null. We defer the restricted Boltzmann machine example to Appendix A.1. The code may be found here.
Figure 1 exhibits the performance of the proposed test with $\alpha=0.05$, under the alternatives $\theta_{1}=1$ and $\theta_{1}=(1,1)$ respectively. We emphasize the exponential growth of the wealth process. This behaviour is expected, in view of Theorem 2 and the fact that all these examples fulfil regularity conditions so that $\mathbb{E}_{H_{1}}[g^{*}(X)]>0$; an actual empirical verification of the lower bound derived in Theorem 2 is deferred to Appendix A.6. Furthermore, we stress that aGRAPA empirically outperforms LBOW. This is due to the fact that $\lambda_{t+1}^{\text{aGRAPA}}>\lambda_{t+1}^{\text{LBOW}}$, so aGRAPA bets more aggressively. We defer the illustration of the type-I error control at the desired level $0.05$ to Appendix A.2, a discussion on how the tightness of the bounds $M_{p}$ empirically affects the power of the test to Appendix A.5, and examples of testing composite null hypotheses to Appendix A.4.
Figure 2: Proportion of rejections for the Gaussian distribution considered in Section 6 under the alternatives $\theta_{1}\in\{0.40,0.42,0.44,0.46,0.48,0.50\}$ for (I) the (classical) batch setting kernelized Stein discrepancy with sample size $n$, (II) the (proposed) sequential kernelized Stein discrepancy. The (proposed) sequential test always ends up rejecting the null hypotheses, while the batch test will not do so if the original sample size is too small.
We emphasize once again that, in contrast to sequential procedures, batch setting algorithms implicitly commit to a sample size that is chosen prior to running the experiment, which may lead to substantially suboptimal choices of sample sizes. We exhibit such a phenomenon in Figure 2, which displays the proportion of rejections for the Gaussian distribution under different alternatives and illustrates the convenience of employing anytime valid tests. If using the classical kernelized Stein discrepancy test (left plot), $20$ observations lead to rejection rates lower than $0.50$, and $400$ observations show rejection rates of $1$. This implies that $20$ observations are not enough to obtain a powerful test, and probably less than $500$ observations would have sufficed to yield high power (resulting in a cheaper experiment). However, this cannot be known beforehand. In stark contrast, all the null hypotheses are eventually rejected by the proposed test (right plot), with the empirical powers being significantly high already after $150$ observations, and always being able to collect more data and keep running the experiment with anytime validity. We defer a longer comparison between the batch and sequential tests to Appendix A.3.
@@ 7 CONCLUSION
We have developed a sequential version of the kernelized Stein discrepancy goodness-of-fit test, which gives way to goodness-of-fit tests that can handle distributions with unknown normalizing constants and can be continuously monitored and adaptively stopped. We have done so by combining tools from testing by betting with RKHS theory, while avoiding assuming uniform boundedness of the Stein reproducing kernel. We have proved the validity of the test, as well as exponential growth of the wealth process under the alternative and mild regularity conditions. Our experiments have exhibited the empirical performance of the test in a variety of scenarios.
@@ Acknowledgements
DMT gratefully acknowledges that the project that gave rise to these results received the support of a fellowship from ‘la Caixa’ Foundation (ID 100010434). The fellowship code is LCF/BQ/EU22/11930075. AR was funded by NSF grant DMS-2310718. The authors would also like to thank the anonymous reviewers for their valuable comments and suggestions.
@@ Appendix A ADDITIONAL EXPERIMENTS
A.1 Logarithmic wealth process of a Gaussian-Bernoulli restricted Boltzmann machine
In Section 6, we displayed the performance of the proposed test for the Gaussian and intractable models. The performance of the proposed test for the remaining model introduced in Section 5, a Gaussian-Bernoulli restricted Boltzmann machine, is now exhibited in Figure 3. In particular, we take $d_{h}=10$, and $d=50$. We sample from it using Gibbs sampling with a burn-in of 1000 iterations. Under the null, we take $b=0$, $c=0$, and matrix $B$ such that $B_{ij}$ is one if visible node $i$ is connected to hidden node $j$, and zero otherwise, with each hidden node connected to five visible nodes. We consider two different alternatives. For one of them, each entry of $B$ is shifted by $0.5$. For the other one, $b=1$.
Figure 3: Average logarithmic wealth alongside $95\%$ empirical confidence intervals for $1000$ simulations under the alternative for (I) the restricted Boltzmann machine with shifted $B$, (II) the restricted Boltzmann machine with bias $b=1$. We highlight once again the exponential growth of the wealth processes.
A.2 Type-I error control
In Section 6, we illustrated the empirical power of the proposed test. We devote this appendix to displaying the empirical type-I error control of the proposed test. Figure 4 exhibits the wealth processes for the examples considered in Section 6 under the null. We highlight that the wealth processes do not cross the threshold $1/0.05$, and hence the nulls are not rejected.
Figure 4: Average wealth alongside $95\%$ empirical confidence intervals for $1000$ simulations under the null for (I) the Gaussian distribution, (II) the intractable model, (III) the restricted Boltzmann machine. We emphasize that the wealth processes do not cross the threshold $1/0.05$, and hence the nulls are not rejected, showing the empirical type-I error control.
We note that Figure 4 exhibits the anytime validity of the tests until round $100$ (the same number of rounds as considered in Section 6). However, we can verify anytime validity without committing to a sample size for those simulations with known normalizing constants. More specifically, let $P$ be the distribution under the null, and $Q$ any other distribution such that $dP/dQ$ is known and we can sample from $Q$. Let $\tau$ denote the random stopping time. Since $\mathbb{P}_{P}(\tau<\infty)=E_{P}[1(\tau<\infty)]=E_{Q}[dP/dQ1(\tau<\infty)]$, instead of sampling from $P$ and calculating $\mathbb{P}_{P}(\tau<\infty)$ by Monte-Carlo, we can instead sample from $Q$ and calculate $E_{Q}[dP/dQ1(\tau<\infty)]$ by Monte-Carlo. For $Q\neq P$, the procedure will stop relatively fast (since the wealth grows exponentially fast under the alternative), and we can average these likelihood ratios at the stopping time to estimate the type 1 error. For simulations in which we know the normalizing constants (even though we may not want to use them in order to test our methods), this importance sampling technique for estimating the type-1 error works well. For the Gaussian case considered in Section 6 (i.e., $P$ is a standard Gaussian distribution), taking $Q$ to be a Gaussian distribution with unit variance centered at $0.5$ leads to an estimated $\mathbb{P}_{P}(\tau<\infty)$ of $0.0006$ for $\alpha=0.1$ (showing, again, the empirical anytime validity of the proposed test).
A.3 Comparison to batch setting
We illustrated in Section 6 some of the advantages of the sequential test over the batch test. We emphasize that the proposed test is anytime valid, counting with the subsequent substantial advantages that have been discussed in the main body of the work. Thus, it is expected to be outperformed in the batch setting by algorithms that are tailored to fixed sample sizes. Nonetheless, it is of interest to explore how the proposed test compares with the more classical kernel Stein discrepancy test. We present in this appendix the empirical performance of both the sequential and fixed sample size algorithms.